# Final-Model-Only Data Attribution with a Unifying View of Gradient-Based Methods

### v1

@@ D.2 Training of Final Models
For all tabular datasets, we used a 2-hidden-layer multi-layer perceptron (MLP) with 128 units in each hidden layer. For the regression problems, the MLP has one output and the loss function is mean squared error (MSE), while for (binary) classification, the MLP has two outputs (logits) and the loss function is cross-entropy. The MLPs were trained using SGD for $T=1000$ epochs and a batch size of 128 (following Bae et al. (2022)) to yield the final model $f({\bm{x}},{\bm{\theta}}^{f})$, except for the larger Folktables dataset where $T=100$. Learning rates were as follows: $0.3$ for Concrete and Energy, $0.001$ for FICO, and $0.01$ for Folktables.
@@
We used a ResNet-9 architecture (He et al., 2015) and trained it using SGD to minimize cross-entropy loss on the CIFAR-10 training set for $T=50$ epochs. We trained with a batch size of 512, learning rate of 0.4, and weight decay of 0.001.
@@
We used a pre-trained bert-base-uncased model [^4] https://huggingface.co/google-bert/bert-base-uncased (Devlin et al., 2019) and attached a linear classification layer that maps the BERT model’s final representation of the CLS token to logits. We fine-tuned this model on the SST-2 training set using AdamW (Loshchilov and Hutter, 2019) to minimize cross-entropy loss, with a batch size of 64, learning rate of $10^{-5}$, zero weight decay, and gradient norm clipping at a threshold of 1. Accuracy on the SST-2 validation set was used to choose the number of epochs from a range of $1$ to $30$. This resulted in a final model with validation accuracy of $92.9\%$ after epoch 19.
@@ D.3 Implementation of further training
For the case in which the further training algorithm $A^{\prime}$ is the same as the initial training algorithm $A$, the learning rate for $A^{\prime}$ was chosen to be one order of magnitude smaller than that for $A$, i.e., $0.03$ for Concrete and Energy, $10^{-4}$ for FICO and Folktables, $0.04$ for CIFAR-10, $10^{-6}$ for SST-2. The reason is that the given parameters ${\bm{\theta}}^{f}$ are closer to convergence than the initial parameters ${\bm{\theta}}^{(0)}$. The maximum number of epochs $T^{\prime}$ for $A^{\prime}$ was also chosen to be a fraction of $T$: $T^{\prime}=500$ for Concrete and Energy following Bae et al. (2022), $T^{\prime}=200$ for FICO, $T^{\prime}=25$ for Folktables, $T^{\prime}=10$ for CIFAR-10, and $T^{\prime}=1$ for SST-2. Other parameters such as batch size were the same as for $A$. For the experiment on tabular data in which $A^{\prime}$ was Adam, smaller learning rates were used since Adam is a more powerful optimizer: $10^{-4}$ for Concrete and Energy, $10^{-6}$ for FICO and Folktables.
@@
Regarding the random seeds (denoted as $\xi^{(s)}$ in (13), (17)) that determine the order of instances in each epoch, this control was achieved in PyTorch (Paszke et al., 2019) as follows: First, the seed of a random number Generator object was set to a specified value (using generator.manual_seed()). Then, a RandomSampler object was instantiated with this Generator (RandomSampler(generator=generator, ...)), and a DataLoader was instantiated with the RandomSampler (DataLoader(sampler=sampler, ...)).
@@
To achieve the effect of leaving out one training instance ${\bm{z}}_{i}$, we followed Schioppa et al. (2023) in subtracting a second loss term $L({\bm{z}}_{i};{\bm{\theta}})/n$ from the usual loss for every batch. The scaling by $1/n$ (recall that $n$ is the number of training instances) approximately cancels out the weight of instance ${\bm{z}}_{i}$ in the usual loss. Denoting by $B$ and $n_{B}=\lceil n/B\rceil$ the batch size and number of batches per epoch, the weight of ${\bm{z}}_{i}$ in the usual loss is $1/B$ (one appearance per epoch), whereas the total weight from the subtracted loss term is $n_{B}/n\approx 1/B$.
@@
Results corresponding to (17) are shown in Figures 5 and 7.
@@ D.4 Approximate TDA methods
We used the implementations of these methods in the torch-influence repository [^5] https://github.com/alstonlo/torch-influence provided by Bae et al. (2022). For both CG and LiSSA, we used a damping value of $\lambda=0.01$, which we found to yield higher cosine similarity values than the other $\lambda=0.001$ value used by Bae et al. (2022). We observed that cosine similarities increase as $\lambda$ increases, so it is possible that further tuning of $\lambda$ may yield better results. For LiSSA’s additional parameters, we set the depth (number of iterations) to 5000, the number of repeats (repeat) to 1, and the scale parameter to the smallest value in $\{10,20,50,100,200,500\}$ for which LiSSA did not diverge.
@@
For the tabular datasets, we used the repository [^6] https://github.com/ykwon0407/DataInf provided by Kwon et al. (2024) with default parameter values, specifically lambda_const_param=10 which determines the damping value. However for CIFAR-10 and SST-2, the authors’ code was not computationally efficient enough. Instead, we implemented DataInf’s key equation (Kwon et al., 2024, eq. (5)) ourselves. In doing so, we approximated the average over all training instances $i=1,\dots,n$ with an average over a random subsample of size $1000$. We also first computed the gradient inner products $L_{l,i}$, $L_{l,ik}$, $L_{l,k}$ in their eq. (5) before computing eq. (5) itself.
@@
We used the repository [^7] https://github.com/pomonam/kronfluence corresponding to Grosse et al. (2023) with default parameter settings.
@@
We used the repository [^8] https://github.com/MadryLab/trak provided by Park et al. (2023) with default parameter settings. For the regression datasets Concrete and Energy, since the TRAK repository does not provide an implementation for regression, we tried to create one ourselves following their documentation. [^9] https://trak.readthedocs.io/en/latest/modeloutput.html. For their “model output function,” we use the model’s prediction, $f({\bm{x}};{\bm{\theta}})$ in our notation. The squared error loss is $L({\bm{z}};{\bm{\theta}})=\frac{1}{2}(f({\bm{x}};{\bm{\theta}})-y)^{2}$, so the output-to-loss gradient that TRAK also requires is
@@ D.5 Evaluation metric
The vectors ${\bm{a}}$ of gold attribution scores and $\hat{{\bm{a}}}$ of approximate attribution scores may have different scales. In this work, we do not concern ourselves with the differing scales, so we take both to be normalized to unit $\ell_{2}$ norm. In this case, the squared Euclidean distance $\lVert\hat{{\bm{v}}}-{\bm{v}}\rVert^{2}$ and cosine similarity $\hat{{\bm{v}}}^{T}{\bm{v}}$ are equivalent up to an affine transformation, and we use the latter as the evaluation metric. Note that cosine similarity is a more demanding measure than Pearson correlation (used for example in Schioppa et al. (2023); Bae et al. (2022)), since the latter corresponds to approximating each gold score $a_{i}$ by $\alpha\hat{a}_{i}+\beta$ where $\beta$ is a non-zero bias, whereas the former constrains $\beta$ to be zero.
D.6 Computing infrastructure
Experiments were run on an internal computing cluster providing nodes with 32 GB of CPU memory, V100 GPUs with 32 GB of GPU memory, and occasionally A100 GPUs with 40 or 80 GB of GPU memory. One CPU and one GPU were used at a time.
@@ Appendix E Additional Experimental Results and Discussion
We reproduce DataInf’s key equation (Kwon et al., 2024, eq. (5)) below (with a sign change to be consistent with Section 4):
@@
where $k$ indexes the training instance being attributed to, $l$ the layers of the NN, and $i$ all training instances. $L_{l,i}$ and $L_{l,ik}$ are inner products between layer-specific loss gradients, as defined in Kwon et al. (2024). The first term in (18) is like Grad-Dot (6) but with a layer-dependent weighting $1/\lambda_{l}$. The remaining average over $i$ can be seen as a correction term. Our results in Figures 25 suggest that this correction term is small.
@@
Different numbers of random seeds averaged
In Figures 6 and 7, we show the counterparts to Figure 3 when further training is done with Adam and when adjustment is done using (17), respectively. As in Figure 3, the cosine similarity increases in all cases with the number of seeds, implying that the averaging of further training brings it closer to what gradient-based methods estimate. The increases are especially notable in Figure 7 when adjustment is done using full-dataset further training (17). As seen in Figure 5, this adjustment method is noisier and appears to benefit more from averaging.
@@
The cosine similarities for CIFAR-10 and SST-2 in Figures 2(e), 2(f) and elsewhere are significantly lower than those for the tabular datasets. As a first step to understanding why, we plot in Figure 8 the gold attribution scores from further training (13), where for each test instance, we have sorted the scores $a_{i}$ in increasing order, and then averaged over the $m$ test instances. We also sort and average the attribution scores from Grad-Dot and Grad-Cos in the same manner. For FICO and Folktables in Figures 8(a), 8(b), the further training attribution scores are more “typical” in that they are approximately symmetric (positive versus negative) and have tails of larger positive and negative values. For CIFAR-10 however (Figure 8(c)), the curves are more “hockey-stick”-shaped, with a heavier positive tail and not much of a negative tail; this pattern becomes more pronounced after 11 steps of further training than after 1 step. In this case, the Grad-Cos curve offers a better match than the Grad-Dot one. For SST-2, the gold attribution score curves are angular: almost all of the scores are close to zero, with only one or two significant non-zero values at either end. In both Figures 8(c) and 8(d), Grad-Dot and Grad-Cos struggle more to approximate these less typical attribution score curves. Further investigation into the poorer approximation quality is ongoing.

### v2

@@ D.2 Training of Final Models
For all tabular datasets, we used a 2-hidden-layer multi-layer perceptron (MLP) with 128 units in each hidden layer. For the regression problems, the MLP has one output and the loss function is mean squared error (MSE), while for (binary) classification, the MLP has two outputs (logits) and the loss function is cross-entropy. The MLPs were trained using SGD for $T=1000$ epochs and a batch size of 128 (following [bae2022infanswer]) to yield the final model $f({\bm{x}},{\bm{\theta}}^{f})$, except for the larger Folktables dataset where $T=100$. Learning rates were as follows: $0.3$ for Concrete and Energy, $0.001$ for FICO, and $0.01$ for Folktables.
@@
We used a ResNet-9 architecture [resnet] and trained it using SGD to minimize cross-entropy loss on the CIFAR-10 training set for $T=50$ epochs. We trained with a batch size of 512, learning rate of 0.4, and weight decay of 0.001.
@@
We used a pre-trained bert-base-uncased model [^4] https://huggingface.co/google-bert/bert-base-uncased [bert] (Apache 2.0 license) and attached a linear classification layer that maps the BERT model’s final representation of the CLS token to logits. We fine-tuned this model on the SST-2 training set using AdamW [loshchilov2018decoupled] to minimize cross-entropy loss, with a batch size of 64, learning rate of $10^{-5}$, zero weight decay, and gradient norm clipping at a threshold of 1. Accuracy on the SST-2 validation set was used to choose the number of epochs from a range of $1$ to $30$. This resulted in a final model with validation accuracy of $92.9\%$ after epoch 19.
@@ D.3 Implementation of further training
For the case in which the further training algorithm $A^{\prime}$ is the same as the initial training algorithm $A$, the learning rate for $A^{\prime}$ was chosen to be one order of magnitude smaller than that for $A$, i.e., $0.03$ for Concrete and Energy, $10^{-4}$ for FICO and Folktables, $0.04$ for CIFAR-10, $10^{-6}$ for SST-2. The reason is that the given parameters ${\bm{\theta}}^{f}$ are closer to convergence than the initial parameters ${\bm{\theta}}^{(0)}$. The maximum number of epochs $T^{\prime}$ for $A^{\prime}$ was also chosen to be a fraction of $T$: $T^{\prime}=500$ for Concrete and Energy following [bae2022infanswer], $T^{\prime}=200$ for FICO, $T^{\prime}=25$ for Folktables, $T^{\prime}=10$ for CIFAR-10, and $T^{\prime}=1$ for SST-2. Other parameters such as batch size were the same as for $A$. For the experiment on tabular data in which $A^{\prime}$ was Adam, smaller learning rates were used since Adam is a more powerful optimizer: $10^{-4}$ for Concrete and Energy, $10^{-6}$ for FICO and Folktables.
@@
Regarding the random seeds (denoted as $\xi^{(s)}$ in (13), (17)) that determine the order of instances in each epoch, this control was achieved in PyTorch [pytorch] as follows: First, the seed of a random number Generator object was set to $s=0,1,\dots,r-1$ ($r=100$) using generator.manual_seed(). Then, a RandomSampler object was instantiated with this Generator (RandomSampler(generator=generator, ...)), and a DataLoader was instantiated with the RandomSampler (DataLoader(sampler=sampler, ...)).
@@
To achieve the effect of leaving out one training instance ${\bm{z}}_{i}$, we followed [schioppa2023theoretical] in subtracting a second loss term $L({\bm{z}}_{i};{\bm{\theta}})/n$ from the usual loss for every batch. The scaling by $1/n$ (recall that $n$ is the number of training instances) approximately cancels out the weight of instance ${\bm{z}}_{i}$ in the usual loss. Denoting by $B$ and $n_{B}=\lceil n/B\rceil$ the batch size and number of batches per epoch, the weight of ${\bm{z}}_{i}$ in the usual loss is $1/B$ (one appearance per epoch), whereas the total weight from the subtracted loss term is $n_{B}/n\approx 1/B$.
@@
Results corresponding to (17) are shown in Figures 5 and 9.
@@ D.4 Approximate TDA methods
We used the implementations of these methods in the torch-influence repository [^5] https://github.com/alstonlo/torch-influence (Apache 2.0 license) provided by [bae2022infanswer]. For both CG and LiSSA, we used a damping value of $\lambda=0.01$, which we found to yield higher cosine similarity values than the other $\lambda=0.001$ value used by [bae2022infanswer]. We observed that cosine similarities increase as $\lambda$ increases, so it is possible that further tuning of $\lambda$ may yield better results. For LiSSA’s additional parameters, we set the depth (number of iterations) to 5000, the number of repeats (repeat) to 1, and the scale parameter to the smallest value in $\{10,20,50,100,200,500\}$ for which LiSSA did not diverge.
@@
For the tabular datasets, we used the repository [^6] https://github.com/ykwon0407/DataInf (license unknown) provided by [kwon2024datainf] with default parameter values, specifically lambda_const_param=10 which determines the damping value. However for CIFAR-10 and SST-2, the authors’ code was not computationally efficient enough. Instead, we implemented DataInf’s key equation [kwon2024datainf, eq. (5)] ourselves. In doing so, we approximated the average over all training instances $i=1,\dots,n$ with an average over a random subsample of size $1000$. We also first computed the gradient inner products $L_{l,i}$, $L_{l,ik}$, $L_{l,k}$ in their eq. (5) before computing eq. (5) itself.
@@
We used the repository [^7] https://github.com/pomonam/kronfluence (Apache 2.0 license) corresponding to [grosse2023studyingllmgen] with default parameter settings.
@@
We used the repository [^8] https://github.com/MadryLab/trak (MIT license) provided by [park2023trak] with default parameter settings. For the regression datasets Concrete and Energy, since the TRAK repository does not provide an implementation for regression, we tried to create one ourselves following their documentation. [^9] https://trak.readthedocs.io/en/latest/modeloutput.html. For their “model output function,” we use the model’s prediction, $f({\bm{x}};{\bm{\theta}})$ in our notation. The squared error loss is $L({\bm{z}};{\bm{\theta}})=\frac{1}{2}(f({\bm{x}};{\bm{\theta}})-y)^{2}$, so the output-to-loss gradient that TRAK also requires is
@@ D.5 Evaluation metric
The derivations in Section 4 indicate that gradient-based methods should approximate the signed magnitudes of further training gold attribution scores. Thus, we chose an evaluation metric that is sensitive to signed magnitudes, not just rankings (evaluated for example by Spearman rank correlation, a common metric in the TDA literature). Having said this, the vectors $\bm{a}$} of gold attribution scores and $\hat{{\bm{a}}}$ of approximate attribution scores may have different scales. In this work, we do not concern ourselves with the differing scales, so we take both to be normalized to unit $\ell_{2}$ norm. In this case, the squared Euclidean distance $\lVert\hat{{\bm{v}}}-{\bm{v}}\rVert^{2}$ and cosine similarity $\hat{{\bm{v}}}^{T}{\bm{v}}$ are equivalent up to an affine transformation, and we use the latter as the evaluation metric. Note that cosine similarity is a more demanding measure than Pearson correlation (used for example in [schioppa2023theoretical, bae2022infanswer]), since the latter corresponds to approximating each gold score $a_{i}$ by $\alpha\hat{a}_{i}+\beta$ where $\beta$ is a non-zero bias, whereas the former constrains $\beta$ to be zero. In addition, we also show results with Spearman correlation in Figure 6, which are qualitatively similar to the cosine similarity results.
D.6 Computing resources
Experiments were run on an internal computing cluster providing nodes with 32 GB of CPU memory, V100 GPUs with 32 GB of GPU memory, and occasionally A100 GPUs with 40 or 80 GB of GPU memory. V100s were sufficient for all training however. One CPU and one GPU were used at a time.
The most computationally expensive part of the experiments was to perform further training on the BERT model for SST-2. Each further training run ($T^{\prime}=1$ epoch) took $\sim 12$ minutes, which includes time to evaluate intermediate checkpoints on the test set after every $\sim 10$ optimization steps. With $l=50$ LOO instances and $r=100$ random seeds, the total time for further training BERT was $\sim 1000$ GPU-hours. The total time for further training ResNet for CIFAR-10 was a significant fraction of this, perhaps $\sim 400$ GPU-hours. Conducting the mislabelled example detection experiment on SST-2 also required $\sim 400$ GPU-hours for further training BERT with $40$ random seeds. Aside from further training, among the gradient-based methods, LiSSA was the most computationally expensive. Running LiSSA and LiSSA-H for CIFAR-10 and SST-2 took $\sim 500$ GPU-hours in total. Thus altogether, a reasonable estimate for total experiment compute is $\sim 3000$ GPU-hours.
@@ Appendix E Additional Experimental Results and Discussion
E.1 Comparison of gradient-based methods to further training
@@
We reproduce DataInf’s key equation [kwon2024datainf, eq. (5)] below (with a sign change to be consistent with Section 4):
@@
where $k$ indexes the training instance being attributed to, $l$ the layers of the NN, and $i$ all training instances. $L_{l,i}$ and $L_{l,ik}$ are inner products between layer-specific loss gradients, as defined in [kwon2024datainf]. The first term in (18) is like Grad-Dot (6) but with a layer-dependent weighting $1/\lambda_{l}$. The remaining average over $i$ can be seen as a correction term. Our results in Figures 25 suggest that this correction term is small.
@@
(e) CIFAR-10
(f) SST-2
Figure 6: Spearman rank correlation between attribution scores of gradient-based TDA methods and further training, as a function of the number of epochs or steps of further training.
Different evaluation metric
In Figure 6, we show a counterpart to Figure 2 where cosine similarity between attribution vectors $\bm{a}$} and $\hat{{\bm{a}}}$ is replaced by Spearman correlation (and then averaged over test instances as before). The plots are qualitatively similar to those in Figure 2, particularly in terms of the grouping of the gradient-based methods. For example in Figure 6(c), DataInf, Grad-Dot, and Grad-Cos remain at the top, while CG and LiSSA are next and again coincide with each other. In Figure 6(d), CG rises and crosses over the top three (DataInf, Grad-Cos, Grad-Dot), similar to Figure 2(d). In Figure 6(f), DataInf and Grad-Dot are again highest and EK-FAC is again in third place.
Different numbers of random seeds averaged
In Figures 7, 8, and 9, we plot the maximum cosine similarity over epochs/steps as a function of the number of random seeds averaged. The detailed procedure is as follows. Given a number $r$ of seeds to be averaged, we divide the 100 seeds that we have in total into $100/r$ groups of $r$ seeds. For each group of $r$ seeds, we obtain a gold attribution vector $\bm{a}$} using either (13) or (17) depending on the adjustment method, and compute its cosine similarity with a gradient-based attribution vector $\hat{{\bm{a}}}$. We then compute means and standard errors of the cosine similarities over the $m$ test instances as well as the $100/r$ groups. This yields mean cosine similarities as a function of further training epochs or steps, like in Figures 25. Finally, we take the maximum over epochs or steps to give us the $y$-axis value in Figures 79 corresponding to the number of seeds $r$ on the $x$-axis. This procedure is repeated for different values of $r$.
In all cases in Figures 79, the cosine similarity increases with the number of seeds, implying that the averaging of further training runs brings it closer to what gradient-based methods estimate. The increases are more notable for the first-order methods and DataInf. This may appear so because the first-order-like methods can achieve higher maximum cosine similarities than second-order methods, provided there is sufficient averaging. The increases are especially notable in Figure 9 when adjustment is done using full-dataset further training (17). As seen in Figure 5, this adjustment method is noisier and appears to benefit more from averaging.
(a) Concrete Strength
(b) Energy Efficiency
(c) FICO
(d) Folktables
(e) CIFAR-10
(f) SST-2
Figure 7: Maximum cosine similarity between attribution scores of gradient-based TDA methods and further training, as a function of the number of random seeds averaged. Here the further training algorithm $A^{\prime}$ is the same as the initial training algorithm.
(a) Concrete Strength
(b) Energy Efficiency
(c) FICO
(d) Folktables
@@
TRAKM=1
There are two possible reasons for the poor approximation quality of TRAKM=1 seen in Figure 2 and elsewhere. First, in the FiMO setting, we have only $M=1$ checkpoint and can only evaluate the TRAKM=1 variant as mentioned in Section 4.3.3, i.e., without ensembling. [park2023trak] show that using an ensemble of $M\gg 1$ checkpoints greatly improves performance. Second, we compute similarity with further training attribution values, whereas [park2023trak] evaluate using their linear datamodelling score (LDS) metric, which is based on re-training and for the TAA setting.
@@
The cosine similarities for CIFAR-10 and SST-2 in Figures 2(e), 2(f) and elsewhere are significantly lower than those for the tabular datasets. As a first step to understanding why, we plot in Figure 10 the gold attribution scores from further training (13), where for each test instance, we have sorted the scores $a_{i}$ in increasing order, and then averaged over the $m$ test instances. We also sort and average the attribution scores from Grad-Dot and Grad-Cos in the same manner. For FICO and Folktables in Figures 10(a), 10(b), the further training attribution scores are more “typical” in that they are approximately symmetric (positive versus negative) and have tails of larger positive and negative values. For CIFAR-10 however (Figure 10(c)), the curves are more “hockey-stick”-shaped, with a heavier positive tail and not much of a negative tail; this pattern becomes more pronounced after 11 steps of further training than after 1 step. In this case, the Grad-Cos curve offers a better match than the Grad-Dot one. For SST-2, the gold attribution score curves are angular: almost all of the scores are close to zero, with only one or two significant non-zero values at either end. In both Figures 10(c) and 10(d), Grad-Dot and Grad-Cos struggle more to approximate these less typical attribution score curves. Further investigation into the poorer approximation quality is ongoing.
Comparison to re-training
Figure 11 shows cosine similarities between attribution scores of gradient-based TDA methods and LOO re-training (from scratch) for the tabular regression datasets. In stark contrast to further training in Figure 2, the gradient-based attribution scores are nearly uncorrelated with LOO re-training.
(a) Concrete Strength
(b) Energy Efficiency
Figure 11: Cosine similarity between attribution scores of gradient-based TDA methods and LOO re-training (from scratch), as a function of the number of epochs of re-training.
E.2 Validation of further training
E.2.1 Mislabelled example detection
Further training is a natural choice for a gold standard because it is the analogue to re-training in the FiMO setting and is approximated by existing gradient-based methods. To further support the validity of this choice, we conducted an experiment to determine how well further training aligns with a downstream application of TDA, namely mislabelled data detection. In the context of FiMO and further training, we pose this as a question of whether the sensitivity of a final model, as measured by further training influence values, is correlated with mislabelled examples. The setup for this experiment is similar to that in Section 6.1 and prior works [kwon2024datainf, koh2017inffunc]. For the binary classification datasets, we flip the labels of a randomly selected 20% of the training subset $\mathcal{L}$ on which we perform LOO further training. We then evaluate the correlation of further training influence values with the indicator of flipped instances, as measured by AUC-ROC following [kwon2024datainf]. After one epoch of further training, the AUC values are as follows (mean and standard error over 40 trials): FICO: $0.758\pm 0.008$; Folktables: $0.627\pm 0.012$; SST-2: $0.995\pm 0.003$. The last value is high because the further training influence values of flipped and non-flipped points are almost perfectly separable. We conclude that further training is indeed aligned with mislabelled data detection.
E.2.2 Qualitative examples
The authors of the gradient-based TDA methods considered in this work have shown that examining the most influential training instances (most positive and negative) for a given test instance can give insights into what the model has learned. We refer the reader to Figure 3 of [kwon2024datainf] and Figure 3 in the arXiv version of [park2023trak] for such examples. [grosse2023studyingllmgen] use such influential examples to make observations about LLM phenomena, such as increasing abstraction with model scale (see their Section 5.3). In the same way that illustrative examples can be provided for approximate TDA methods, we can do the same for the further training gold standard. However, given the computational cost of further training and the experimental setup in Section 6.1, we are limited to identifying most influential examples within the evaluated subset $\mathcal{L}$} of training instances. With this limitation in mind, we discuss our findings for two of the datasets.
Folktables
In Table 2, we show the training instances with the most positive and most negative influence values for the first two test instances evaluated from the Folktables dataset. Since the most positive influence value implies that the loss on the test instance increases when the training instance is left out, we call this the most “helpful” training instance, and similarly call the training instance with the most negative influence value the most “harmful”. Similar to previous work (for example Figure 3 in the arXiv version of [park2023trak]), the most influential training instances are semantically similar to the test instance, and the most helpful instances tend to have the same label while the most harmful have the opposite label. For example, the first test instance is a 61-year-old white man who likely has a blue-collar profession (“Sailors, Marine Oilers, Ship Engineers”) and earns more than 50,000 USD a year (high income). The most helpful and harmful training instances are also white men with blue-collar jobs (construction laborer and structural metal fabricator/fitter). The most helpful instance also earns more than 50,000 and is younger at 38 years old, while the most harmful instance earns less than 50,000 and is older at 84.
Table 2: Most influential training instances for the first two evaluated test instances from the Folktables dataset.
Feature Test instance Most helpful training instance Most harmful training instance
Label (income) High High Low
Age 61 38 84
Work hours 42 15 29
Category of worker Employee of private for-profit Employee of private for-profit Employee of private for-profit
Schooling Associate’s degree Master’s degree No schooling completed
Marital status Married Never married Married
Occupation Sailors, Marine Oilers, Ship Engineers Construction Laborers Structural Metal Fabricators & Fitters
Place of birth Massachusetts New Jersey Massachusetts
Sex Male Male Male
Race White alone White alone White alone
Label (income) High High Low
Age 39 24 56
Work hours 50 60 45
Category of worker Employee of private for-profit Employee of private for-profit Employee of private for-profit
Schooling GED or alternative credential Bachelor’s degree Some college, << 1 year
Marital status Married Never married Married
Occupation Office Clerks, General 1st-Line Supervisors Of Food Workers Lic. Practical & Vocational Nurses
Place of birth Massachusetts Connecticut Massachusetts
Sex Female Male Female
Race White alone White alone White alone
SST-2
We also examined the most influential training instances within $\mathcal{L}$} for the $m$ test instances that we evaluated from SST-2, both individually as well as an average over the test instances. Interestingly, we find that in almost all cases, the same two training instances are most influential. These are the texts “outnumber the hits by three-to-one” (negative sentiment) and “a fresh infusion” (positive sentiment). We thus infer that these two instances were the most important within $\mathcal{L}$} for the BERT model to learn in order to classify sentiment well in general. We also note that “outnumber the hits by three-to-one” is a more difficult example to classify as negative, since it requires understanding “hits” (normally positive) and “outnumber” and combining the two.
Appendix F Additional Limitations and Future Work
Below we summarize some additional limitations of this work that were mentioned in passing in the main text:
  • FiMO setting: As mentioned in Section 3, a limitation of the FiMO setting is that it does not seem to permit estimation of the contribution of a training instance to the final model by “going back in time.” It may only be possible to measure sensitivity to a training instance (i.e., a localized notion of contribution). In addition, a reviewer noted that sensitivity may be difficult to measure if the model has saturated on a training instance.
  • Approximation by gradient-based methods: The fundamental assumption for gradient-based methods to approximate further training is that the amount of further training is limited, as stated at the beginning of Section 4. This limitation is evidenced by the decay of the cosine similarity curves in Figure 2 and elsewhere, especially of the first-order methods.
We elaborate on the additional points for future work listed in Section 7.
  1. 1)
    An open question is how much further training is the right amount to determine the final model’s sensitivity to training instances. If too little, the sensitivity may not be measurable, whereas if too much, then the measurement may no longer be local, i.e., specific to the given model $f({\bm{x}};{\bm{\theta}}^{f})$. The extent to which further training is well-approximated by Taylor expansions (as evaluated for example in Figure 2) may provide a partial answer.
  2. 2)
    We restricted attention in this work to LOO further training, as mentioned in Section 3. Future research on TDA in general could focus more on units larger than a single instance (referred to as “group influence” [koh2019group, basu2020second, hu2024most]), beyond making the additivity assumption that the attribution score for a group is the sum of LOO attribution scores.
  3. 3)
    While we have presented further training as applied to the original training set $\mathcal{D}$ and assumed that $\mathcal{D}$ is available, there is no obstacle to considering further training on new, unseen instances. This would broaden the problem to predicting the effect of further training on new data and connect to problems such as data selection.
Appendix G Broader Impacts
This paper highlights the final-model-only setting for TDA, which is a setting that allows a wider set of practitioners (for example HuggingFace model users) to perform TDA because it does not assume access to the training algorithm or intermediate information from training. At the same time, this wider access to TDA might lead to increased misinterpretation or misplaced trust in attribution scores, especially among less expert users. Modern neural networks and neural network training procedures are highly complex, and any insights obtained into them should be tempered with some scientific skepticism. Beyond these considerations, there may be other potential societal consequences of our work, in line with machine learning research in general. Hence we do not feel that they must be specifically highlighted here.